How would you prepare 425 g of an aqueous solution containing 2.40% by mass of sodium acetate, NaC2H3O2?
Answer #1

425 g aqueous solution = sodium acetate + water

According to the question, sodium acetate present 2.40 % by mass. It means 100 g aqueous solution contain 2.40 g of sodium acetate and 97.60 g of water.

Since 100 g solution contain = 2.40 g of sodium acetate

1 g solution contain = 2.40 g/100 = 0.0240 g

425 g solution contain = 0.0240 g × 425 = 10.2 g of sodium acetate

Therefore mass of water required = 425 g - 10.2 g = 414.8 g

Conclusion : 425 g of aqueous solution prepared by dissolving 10.2 g of sodium acetate in 414.8 g of water.